OpenStudy (cherryfrost27):

the sides of a triangle are in the ratio 3:4:5.If the longest side of the triangle is 40 in longer than the shortest side,find the preimeter of the triangle

OpenStudy (cherryfrost27):

@AloneS

jhonyy9 (jhonyy9):

cherry than you check these numbers 3,4 and 5 what triplet are these ?

jhonyy9 (jhonyy9):

3^2 +4^2 =5^2 hope this will help you understanding my question

OpenStudy (cherryfrost27):

so 9 +16 = 25?

jhonyy9 (jhonyy9):

yes but these has the name of pythagorean triplet or trio

jhonyy9 (jhonyy9):

do you know the pythagora"s theorem ?

OpenStudy (cherryfrost27):

no...

jhonyy9 (jhonyy9):

@AloneS please how can you explain to asker than that in this case this is a right triangle than she or he dont learned again about pythagoras theorem ?

OpenStudy (cherryfrost27):

i dont want the answer but do you know the answer to this question?

jhonyy9 (jhonyy9):

yes but i like help you understanding it clearly easy

OpenStudy (cherryfrost27):

yeah thanks! i really need to understand on how to do this

jhonyy9 (jhonyy9):

but than you dont know that from this ratio of 3,4,5 result that this triangle is a right triangle so idk. how can i explain this to you ?

OpenStudy (cherryfrost27):

idk ....

OpenStudy (benji):

The pythagorean theorem is : If a triangle has sides ABC where C is the hypotenuse, then \[A^2+B^2=C^2\]

OpenStudy (cherryfrost27):

oooooooooooooo

jhonyy9 (jhonyy9):

Benji you talk about a right triangle ,yes ? please clearify

jhonyy9 (jhonyy9):

just bc. than your triangle not is right so this equation not is true

OpenStudy (cherryfrost27):

wait so what do i need to do first with this question?

jhonyy9 (jhonyy9):

Benji this is your way - please help

jhonyy9 (jhonyy9):

cherry do you know what mean hypotenuse ?

OpenStudy (cherryfrost27):

the longest side of a right triangle, opposite the right angle.

jhonyy9 (jhonyy9):

exactly very nice

OpenStudy (benji):

This is what we know so far. The Longest side (hypotenuse) is 40inches longer than the shortest side (x).

OpenStudy (cherryfrost27):

ok

jhonyy9 (jhonyy9):

GREAT Benji

OpenStudy (cherryfrost27):

we are still trying to find the perimeter right?

OpenStudy (cherryfrost27):

?

jhonyy9 (jhonyy9):

Benji you not can continuing ?

OpenStudy (cherryfrost27):

benji help please

jhonyy9 (jhonyy9):

use this given ratio of sides 3,4 and 5

jhonyy9 (jhonyy9):

Benji you know that the hypothenuse has a length of x+40 and the shorter side has a length of x how you wrote above use pythagorean theorem and calculi the length of second side

OpenStudy (benji):

Sorry I had to figure it out real quick, ok so we have a ratio of 3:4:5, that means no matter what the lengths of the sides are, they can all be divided back down to 3:4:5, okay? So we have to find a number that can be divided by 3 and a number that can be divided by 5, that is also 40 more than the number that can be divided by 3.

OpenStudy (benji):

So if we start with 3:4:5 and multiply them all by say 10, we have 30:40:50, okay? The largest length is only 20 more than the smallest, so we need bigger numbers.

OpenStudy (benji):

Cherry, pick another number greater than 10 and let's see if we get it.

OpenStudy (cherryfrost27):

okay um 15

OpenStudy (benji):

Alright, 15 would give us 45:60:75, does that work?

OpenStudy (cherryfrost27):

yes?

OpenStudy (cherryfrost27):

wait but

OpenStudy (benji):

We need the difference to be 40, the difference of 75 and 45 is only 30, so we need a bigger number.

OpenStudy (cherryfrost27):

20

OpenStudy (cherryfrost27):

lets try 20 then

OpenStudy (cherryfrost27):

its 20

OpenStudy (benji):

Ok, 20 gives us 60:80:100, that's right! So we have:

OpenStudy (benji):

Now we need the perimeter, which is just the sum of all the sides-as if we were to walk around the whole triangle, what is the perimeter?

OpenStudy (cherryfrost27):

240

OpenStudy (cherryfrost27):

!

OpenStudy (benji):

Right

OpenStudy (cherryfrost27):

so 240 is the answer

OpenStudy (benji):

240 inches is the answer yes

OpenStudy (cherryfrost27):

OMG THANK YOU SO MUCH! <3

jhonyy9 (jhonyy9):

Benji you are a very smart helper - perfect explained - very very nice job GREAT really

OpenStudy (benji):

Yep :)

OpenStudy (benji):

ty @jhonyy9

OpenStudy (cherryfrost27):

you guys are the best thank you so much :)

jhonyy9 (jhonyy9):

np welcome and hope i will can collaborate with you more much in future

jhonyy9 (jhonyy9):

good luck cherry hope you ve understood it all right easy every steps

OpenStudy (cherryfrost27):

thanks !

jhonyy9 (jhonyy9):

np welcome bye bye and anytime just tag me

OpenStudy (cherryfrost27):

alright :)